# An error-mitigated photonic quantum circuit Born machine

### v2

@@ Abstract
Generative machine learning models aim to learn the underlying distribution of the data in order to generate new samples. Quantum circuit Born machines (QCBMs) are a popular choice of quantum generative models which can be implemented on shallow circuits. Within the framework of photonic quantum computing, we present a QCBM designed for linear optics. We show that a new error mitigation technique, called recycling mitigation, greatly improves the training of QCBMs in realistic scenarios with photon loss, both through simulations and an experiment on a quantum photonic integrated processor.
@@ I Introduction
In this work, after introducing a QCBM scheme for linear optical circuits, we show that its training is significantly improved by the recycling mitigation technique in realistic scenarios with photon loss. We demonstrate several situations where a QCBM becomes trainable if recycling mitigation is applied, including a proof-of-concept experiment on our photonic quantum processor Altair. In the discussion section, we consider the question of classical simulability, hardness, and the link with boson sampling [30].
@@ II A photonic QCBM
At the output of the circuit, the state is measured by photon detectors. Different Fock states |$\ket{\vec{n}_{out}}=\ket{n^{out}_{1},...,n^{out}_{m}}$ are detected as arrangements of photons in the output modes, that we denote $s=(n^{out}_{1},...,n^{out}_{m})$. With photon loss, the $n^{out}_{i}$ may not sum to $n$. Detecting all possible output states is not straightforward in practice because photon number resolving detectors are not readily available with current technology. To match experimental settings, we make the simplifying choice of using threshold detectors in most of our simulations, thus only allowing $n^{out}_{i}$ to be 00 or $1$ (‘click’ or ‘no click’). We refer to the Appendix for a note on no-collision regimes and recycling mitigation.
@@
Training strategy. The model is trained as a variational algorithm, which means that a classical optimization procedure is applied to the model parameters $\vec{\theta}$, as shown in Figure 1. In most simulations, we use the Kullback-Leibler (KL) divergence [34] as a loss function. We recall the definition of the KL divergence, given two probability distributions $Q$ and $P$ defined on the same sample space $\mathcal{X}$:
$D_{\text{KL}}(P||Q)=\sum_{x\in\mathcal{X}}P(x)\text{log}\left(\frac{P(x)}{Q(x)}\right).$ (1)
We also consider as metrics the Total Variational Distance (TVD) and the Maximum Mean Discrepancy (MMD) [35]. We refer the reader to the last section of this paper for a discussion around the choice of the loss function, and to the Appendix for more details on model training.
@@ III Recycling mitigation
where $0\leq p_{1}\leq 1$, $p_{id}(s)$ is the ideal probability of outcome $s$ in the absence of photon loss, and $I(s)$ is the interference term, which is an artifact of constructing recycled probabilities from probabilities affected by photon loss. Applying certain classical post-processing techniques, detailed in [29], to the set of recycled probabilities $\{p_{R}(s)\}$ allows the recovery of a set of mitigated probabilities $\{p_{mit}(s)\}$, which are approximations of the ideal probabilities $\{p_{id}(s)\}$.
Recycled probabilities are constructed in such a way as to maximise the weight $p_{1}$, corresponding to the signal of the ideal probability. Furthermore, expectation values of the interference term $\mathbb{E}_{s}(I(s))$ can be computed exactly, this allows the computation of concentration inequalities for $I(s)$. These facts, combined with the fact that the lossy probabilities used in constructing recycled probabilities have smaller statistical errors than post-selected probabilities, are central to why recycling mitigation outperforms post-selection up to large sample sizes.
@@ IV Numerical simulations
We observe that for given $N$ and $\eta$, recycling mitigation significantly improves the training of lossy QCBMs and brings the value of the loss function close to the lossless one, while the lossy unmitigated case remains mostly untrainable. We also note the importance of initialization parameters and the variance in the training process in Figure 2(b). Note that simulation results are displayed for loss parameter $\eta=0.8$, which is within reach for near-term hardware [21]. Our results remain valid for higher loss regimes and values of $\eta$, as we show in the next section.
@@ V Experiment
For our experimental setup, we use the Altair quantum photonic integrated processor. We implement the same scenario as in Figure 2(a), although the main difference is that we find ourselves in a high loss regime: $\eta$ is estimated to be around 0.96. Additionally, the device is affected by other types of noise, such as photon distinguishability, which would require a different kind of error mitigation. Further technical details on the setup can be found in the AppendixC.
@@ VI Discussion
Nevertheless, this particular approach to training retains potential, as it turns out that there is still a polynomial advantage in estimating probabilities given by the modulus squared of a permanent in a boson sampling experiment. Indeed, we show in the Appendix that statistics collected from $t$ i.i.d runs of a boson sampler allow, with high probability, for estimating the modulus squared of the permanent of an arbitrary matrix to a smaller additive error, as compared to the additive error estimate obtained from Gurvits algorithm with a run time of $O(t)$. While these results hold for the lossless case only, this leads us to the open question: could recycling mitigation make these claims robust to noise? Finally, with respect to the recent results of [47], we point out that they do not apply to our setup as we do not consider expectation values of Pauli operators, but a linear optical evolution.
In future work, it would be interesting to extend recycling mitigation techniques to work with implicit loss functions like the MMD, or the Sinkhorn divergence studied in [43]. These would allow for the training phase to be based on samples only, which would also be key to scaling the model. Other future research directions include developing mitigation techniques for other types of photonic errors, as well as exploring optimization in the context of photonics, focusing on topics such as gradient evaluation or parameter initialization.
@@ VIII Acknowledgements
The authors would like to thank Eric Bertasi and Marion Fabre for their precious help improving the code, as well as Joseph Bowles and Nicolas Maring for fruitful discussions and feedback on the manuscript. This work has been co-funded by the European Commission as part of the EIC accelerator program under the grant agreement 190188855 for SEPOQC project, by the Horizon-CL4 program under the grant agreement 101135288 for EPIQUE project, and by the OECQ Project financed by the French State as part of France 2030.
Appendix A Recycling mitigation and no-collision regime
@@
Appendix B Model training and hyperparameters
@@
Gradient evaluation in quantum machine learning is less straightforward than in classical machine learning, due to the lack of an equivalent for the backpropagation algorithm. The parameter-shift method [49, 50] is the current standard for variational quantum circuits – and only very recent work has focused on photonic quantum circuits [51, 52, 53]. Here, we use an optimizer based on the Simultaneous Perturbation Stochastic Approximation (SPSA) method [54]. It involves stochastic gradient approximation, thus allowing for fewer evaluations of the model at each step, which makes it practical for experiments. We start the optimization by randomly initializing the circuit parameters $\vec{\theta}$.
@@
During our training process, we compared different ansätze, or hyperparameters. We explored the case where $k>1$, referring to Figure 1(a) of the main text, including the case where the parameters are repeated between blocks $\vec{\theta_{1}}=\vec{\theta_{2}}=...=\vec{\theta_{k}}$. We also tested the case where each Mach-Zender interferometer in a block is parametrized by only one phase, the other one being set to 0, thus halving the total number of parameters. Overall, we observed the best results for a simple ansatz made of a single block ($k=1$), keeping two parameters per interferometer. This scenario corresponds to the results displayed in Figure 2 of the main text.
Appendix C Experimental setup
@@
The Altair processor presents a similar architecture to the one presented in [21]. It consists of a single-photon source, a 10-mode active demultiplexer followed by fiber delays, a 20-mode photonic integrated circuit, and superconducting nanowire single-photon detectors (SNSPD). The single-photon source is a gated InGaAs quantum dot embedded inside a micropillar cavity. The quantum dot is excited by a laser to produce single photons on demand. At the time of the experiment, single-photon purity was estimated around $0.025\pm 0.002$ and photon indistinguishability around $0.84\pm 0.03$. The train of collected photons is then sent into the demultiplexer, which converts it into up to $n$ photons arriving simultaneously into different modes of the chip. The architecture of the chip is a universal interferometer (as in Figure 1(b) of the main text). It contains 380 phases which are controlled and tuned thermo-optically. The SNSPD detectors are threshold detectors, which means that they can detect the presence of photons (‘click’ or ‘no click’) and not the number of photons in each mode. The detection events are recorded via a time-to-digital converter. The Altair processor is connected to a cloud platform [22], which allows specific operations to be sent to the processor, and which returns the results as detection statistics.
Appendix D Approximating output probabilitities
@@
$\big{|}|\mathsf{E}(\mathsf{Per}(A))|-|\mathsf{Per}(A)|\big{|}\leq|\mathsf{E}(\mathsf{Per}(A))-\mathsf{Per}(A)|\\ \leq\epsilon||A||^{n},$
@@
$|\mathsf{E}(\mathsf{Per}^{2}(A))-\mathsf{Per}^{2}(A)|\leq\epsilon||A||^{n}(2|\mathsf{Per}(A)|+\epsilon||A||^{n})\\ \leq\epsilon(2+\epsilon)||A||^{2n},$
@@
$\big{|}|\mathsf{E}(\mathsf{Per}^{2}(A))|-|\mathsf{Per}^{2}(A)|\big{|}\leq|\mathsf{E}(\mathsf{Per}^{2}(A))-\mathsf{Per}^{2}(A)|\\ \leq\epsilon(2+\epsilon)||A||^{2n},$

### v3

# Error-mitigated photonic quantum circuit Born machine

@@ Abstract
In this article, we study quantum circuit Born machines (QCBMs) in the context of photonic quantum computing. QCBMs are a popular choice of quantum generative machine learning models, and we present a QCBM designed for linear optics. We show that a recently developed error mitigation technique called recycling mitigation greatly improves the training of QCBMs in realistic scenarios with photon loss, which is the primary source of noise in photonic systems. We demonstrate this through numerical simulations and through an experiment on a quantum photonic integrated processor. We expect our work to pave the way towards more demonstrations of error mitigation techniques tailored to photonic devices which can enhance the performance of a quantum algorithm.
@@ I Introduction
In this work, after introducing a QCBM scheme for linear optical circuits, we show that its training is significantly improved by the recycling mitigation technique in realistic scenarios with photon loss. We demonstrate several situations where a QCBM becomes trainable if recycling mitigation is applied, including a proof-of-concept experiment on our photonic quantum processor Altair. In the discussion, we consider the question of classical simulability, hardness, and the link with boson sampling [30, 31].
@@ II A photonic QCBM
There are $m(m-1)/2$ Mach-Zender elements per universal interferometer as depicted in Figure 1(b), and thus $m(m-1)$ phases or parameters per block. While a sequence of unitaries can always be rewritten as one single unitary and thus implemented with one block, we consider it interesting to explore whether this parametrization structure can impact training performance, along with other parametrization choices such as setting all $\theta^{j}_{i}$ to a fixed value and optimizing only over the $\theta^{\prime j}_{i}$. In practice however, we obtained the best numerical results for $k=1$.
At the output of the circuit, the state is measured by photon detectors. Different Fock states |$\ket{\vec{n}_{out}}=\ket{n^{out}_{1},...,n^{out}_{m}}$ are detected as arrangements of photons in the output modes, that we denote $s=(n^{out}_{1},...,n^{out}_{m})$. With photon loss, the $n^{out}_{i}$ may not sum to $n$. Detecting all possible output states is not straightforward in practice because photon number resolving detectors are not readily available with current technology. To match experimental settings, we make the simplifying choice of using threshold detectors in most of our simulations, thus only allowing $n^{out}_{i}$ to be 00 or $1$ (‘click’ or ‘no click’). We refer to the Supplemental Material [35] for more details on no-collision regimes and recycling mitigation.
@@
Training strategy. The model is trained as a variational algorithm, which means that a classical optimization procedure is applied to the model parameters $\vec{\theta}$, as shown in Figure 1. In many of our simulations, we use the Kullback-Leibler (KL) divergence [36] as a loss function. We recall the definition of the KL divergence, given two probability distributions $Q$ and $P$ defined on the same sample space $\mathcal{X}$:
$KL(P||Q)=\sum_{x\in\mathcal{X}}P(x)\text{log}\left(\frac{P(x)}{Q(x)}\right).$ (1)
We also use the Total Variational Distance (TVD) which can be written on discrete domains as [37]:
$\text{TVD}(P,Q)=\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-Q(x)|.$ (2)
Both the KL and the TVD measure how different $Q$ is from $P$. The lower the value of the metric, the more similar the distributions are, and a value of 0 can reached if the two distributions are identical. This is not always desirable as it may indicate overfitting or memorization of the model, and slightly higher values could imply better generalization properties as pointed out in [38].
In generative learning, additional metrics are often used outside of the loss function to evaluate new model samples, sometimes with respect to a test set. For the generation of images, examples include the Fréchet inception distance [39] or the structural similarity index measure [40]. We consider this to be outside the scope of this article, as we focus on the impact of recycling mitigation on loss function behaviour.
Other popular choices of loss functions for quantum models include the Maximum Mean Discrepancy (MMD) [41] or the Sinkhorn divergence studied in [42]. We refer to the last section of this paper for a discussion around the choice of the loss function, and to the Supplemental Material [35] for more details on model training.
@@ III Recycling mitigation
where $0\leq p_{1}\leq 1$, $p_{id}(s)$ is the ideal probability of outcome $s$ in the absence of photon loss, and $I(s)$ is the interference term, which is an artifact of constructing recycled probabilities from probabilities affected by photon loss. Recycling mitigation consists of applying certain classical post-processing techniques, detailed in [29], to the set of recycled probabilities $\{p_{R}(s)\}$, resulting in a set of mitigated probabilities $\{p_{mit}(s)\}$, which are approximations of the ideal probabilities $\{p_{id}(s)\}$.
On an intuitive level, recycling mitigation uses lossy photon statistics in order to get estimates of the ideal (lossless) probabilities. Crucially, the goal of recycling mitigation is to beat post-selection, in the sense of getting lower statistical errors on the estimated ideal probabilities. Post-selection is an unbiased estimator of the ideal probability with the only error being a statistical error $\epsilon_{stat,post}$. Conversely, recycling mitigation results in a biased estimator of the ideal probability, with a bias error $\epsilon_{bias}$ as well as a statistical error $\epsilon_{stat,rec}$. Because recycling mitigation uses the (more probable) lossy statistics to construct its estimators, then $\epsilon_{stat,rec}<\epsilon_{stat,post}$ by construction. Furthermore, it was shown in [29] that the bias error is small enough such that the inequality
$\epsilon_{stat,rec}+\epsilon_{bias}\leq\epsilon_{stat,post},$ (4)
holds for up to a very large number of samples, and for a generic choice of interferometer. This makes recycling mitigation a powerful error mitigation technique applicable to near-term photonic quantum hardware.
@@ IV Numerical simulations
We observe that for given $N$ and $\eta$, recycling mitigation significantly improves the training of lossy QCBMs and brings the value of the loss function close to the lossless one, while the lossy unmitigated case remains mostly untrainable. While the steadily decreasing trend of the loss function shows that the model is learning, we note that there is room for improvement in the model optimization, which would be interesting to study in future work. In particular, the importance of initialization parameters and the variance in the training process can be observed in Figure 2(b). Note that simulation results are displayed for loss parameter $\eta=0.8$, which is within reach for near-term hardware [21]. Our results remain valid for higher loss regimes and values of $\eta$, as we show in the next section.
@@ V Experiment
For our experimental setup, we use the Altair quantum photonic integrated processor. We implement the same scenario as in Figure 2(a), although the main difference is that we find ourselves in a high loss regime: $\eta$ is estimated to be around 0.96. Additionally, the device is affected by other types of noise, such as photon distinguishability [45, 46], which would require a different kind of error mitigation [47]. Further technical details on the setup can be found in the Supplemental Material [35].
@@ VI Discussion
Nevertheless, this particular approach to training retains potential, as it turns out that there is still a polynomial advantage in estimating probabilities given by the modulus squared of a permanent in a boson sampling experiment. Indeed, we show in the Supplemental Material [35] that statistics collected from $t$ i.i.d runs of a boson sampler allow, with high probability, for estimating the modulus squared of the permanent of an arbitrary matrix to a smaller additive error, as compared to the additive error estimate obtained from Gurvits algorithm with a run time of $O(t)$. While these results hold for the lossless case only, this leads us to the open question: could recycling mitigation make these claims robust to noise? Finally, with respect to the recent results of [56], we point out that they do not apply to our setup as we do not consider expectation values of Pauli operators, but a linear optical evolution.
In future work, it would be interesting to extend recycling mitigation techniques to work with implicit loss functions like the MMD or the Sinkhorn divergence. These would allow for the training phase to be based on samples only, which would also be key to scaling the model [55]. Other future research directions include developing mitigation techniques for other types of photonic errors, as well as improving model optimization in the context of photonics, focusing on topics such as gradient evaluation or parameter initialization, or considering circuit components beyond linear optics such as adaptivity and feedforward.
@@ VIII Acknowledgements
The authors would like to thank Eric Bertasi and Marion Fabre for their precious help improving the code, as well as Joseph Bowles and Nicolas Maring for fruitful discussions and feedback on the manuscript. This work has been co-funded by the European Commission as part of the EIC accelerator program under the grant agreement 190188855 for the SEPOQC project; by the Horizon-CL4 program under the grant agreement 101135288 for the EPIQUE project; and by the QuantERA project ResourceQ under Grant Agreement ANR-24-QUA2-007-003.
Appendix A Supplemental Material
Appendix B Recycling mitigation and no-collision regime
@@
Appendix C Model training and hyperparameters
@@
Gradient evaluation in quantum machine learning is less straightforward than in classical machine learning, due to the lack of an equivalent for the backpropagation algorithm. The parameter-shift method [58, 59] is the current standard for variational quantum circuits – and only very recent work has focused on photonic quantum circuits [60-62]. Here, we use an optimizer based on the Simultaneous Perturbation Stochastic Approximation (SPSA) method [63]. It involves stochastic gradient approximation, thus allowing for fewer evaluations of the model at each step, which makes it practical for experiments. We start the optimization by randomly initializing the circuit parameters $\vec{\theta}$.
@@
During our training process, we compared different ansätze, or hyperparameters. We explored the case where $k>1$, referring to Figure 1a of the main text, including the case where the parameters are repeated between blocks $\vec{\theta_{1}}=\vec{\theta_{2}}=...=\vec{\theta_{k}}$. Interestingly, the repetition of gates has been shown to be useful for model properties in certain contexts [64]. We also tested the case where each Mach-Zender element between two modes in a block is parametrized by only one phase ($\theta^{\prime j}_{i}$ in Figure 1 of the main text), while the other one ($\theta^{j}_{i}$) is set to 0, thus halving the total number of parameters. Overall, we observed the best results for a simple ansatz made of a single block ($k=1$), keeping both parameters per Mach-Zender element. This scenario corresponds to the results displayed in Figure 2 of the main text. As mentioned in the main text, a sequence of unitaries can be rewritten as a single unitary and thus implemented with one universal interferometer only. In that case, the choice $k>1$ would correspond to a scenario where each phase in the circuit would be determined by more than one parameter of the model.
Appendix D Experimental setup
@@
The Altair processor presents a similar architecture to the one presented in [21]. It consists of a single-photon source, a 10-mode active demultiplexer followed by fiber delays, a 20-mode photonic integrated circuit, and superconducting nanowire single-photon detectors (SNSPD). The single-photon source is a gated InGaAs quantum dot embedded inside a micropillar cavity. The quantum dot is excited by a laser to produce single photons on demand. At the time of the experiment, single-photon purity was estimated around $0.025\pm 0.002$ and photon indistinguishability around $0.84\pm 0.03$. The train of collected photons is then sent into the demultiplexer, which converts it into up to $n$ photons arriving simultaneously into different modes of the chip. The architecture of the chip is a universal interferometer (as in Figure 1b of the main text). Since the chip has $20$ modes, it contains $20\times 19=380$ phases which are controlled and tuned thermo-optically. Note that a $12$-mode experiment can easily be implemented on a $20$-mode chip by injecting photons only into chosen modes among the first $12$, and blocking transmission into the last $8$ by controlling the phases. The SNSPD detectors are threshold detectors, which means that they can detect the presence of photons (‘click’ or ‘no click’) and not the number of photons in each mode. The detection events are recorded via a time-to-digital converter. The Altair processor is connected to a cloud platform [22], which allows specific operations to be sent to the processor, and which returns the results as detection statistics.
Appendix E Approximating output probabilitities
@@
$\big{|}|\mathsf{E}(\mathsf{Per}(A))|-|\mathsf{Per}(A)|\big{|}\leq|\mathsf{E}(\mathsf{Per}(A))-\mathsf{Per}(A)|\leq\epsilon||A||^{n},$
@@
$|\mathsf{E}(\mathsf{Per}^{2}(A))-\mathsf{Per}^{2}(A)|\leq\epsilon||A||^{n}(2|\mathsf{Per}(A)|+\epsilon||A||^{n})\leq\epsilon(2+\epsilon)||A||^{2n},$
@@
$\big{|}|\mathsf{E}(\mathsf{Per}^{2}(A))|-|\mathsf{Per}^{2}(A)|\big{|}\leq|\mathsf{E}(\mathsf{Per}^{2}(A))-\mathsf{Per}^{2}(A)|\leq\epsilon(2+\epsilon)||A||^{2n},$
@@
We note that, as $n$ grows, the probability encoding $\mathsf{Per}(A)$, i.e $p=\frac{|\mathsf{Per}(A)|^{2}}{||A||^{2n}}$ decreases exponentially with $n$. This can be seen by noting that $\frac{|\mathsf{Per}(A)|}{||A||^{n}}$ is typically upper bounded by a function which is decreasing as $e^{-O(n)}$ [31]. This is consistent with the fact that the size of the output space of the boson sampler grows exponentially with increasing $n$. In the context of Theorem 2, this would mean that, in order to asymptotically get accurate estimates of $p$, $\epsilon$ must decrease exponentially with $n$, and consequently the number of samples collected $O(\frac{1}{\epsilon^{2}})$ must increase exponentially with $n$. We note that this conclusion about the $\epsilon$ scaling also applies to Theorem 1.